L-1{ e-2s / s3}


1
Expert's answer
2021-09-07T00:51:18-0400
"L^{-1}\{\dfrac{e^{-2s}}{s^3}\}=H(t-2)L^{-1}\{\dfrac{1}{s^3}\}(t-2)"

"=H(t-2)\dfrac{(t-2)^2}{2}=\dfrac{H(t-2)\cdot(t-2)^2}{2}"